OpenStudy (anonymous):

A baseball player has been improving every season since making it to the big leagues. Below is a table of the runs he has scored. His manager wants to try to determine when he will score 256 runs. Season Runs 1 2 2 4 3 8 Explain how to create the exponential function that represents his run-scoring abilities. Then explain how to convert this function into a logarithmic function and why this can help the manager answer his question.

OpenStudy (anonymous):

@Deivyneee @CrazyCountryGirl @inkyvoyd @KlOwNlOvE @Lyssa123 @ajprincess @sammixboo @William.deems

OpenStudy (anonymous):

As you can see 1 match = 2runs =2^1 2match= 4 runs = 2^2 2 match = 8 runs= 2^3 so its like 2^n where n= no. of matches

OpenStudy (anonymous):

oh ok thanks so much! @Lyssa123